Solve (𝐷


2 βˆ’ 3𝐷 + 2)𝑦 = π‘₯


2 + sin π‘₯

1
Expert's answer
2021-12-30T05:41:45-0500

characteristic equation:

"k^2-3k+2=0"

"k=\frac{3\pm \sqrt{9-8}}{2}"

"k_1=1,k_2=2"

complementary solution:

"y_c=c_1e^x+c_2e^{2x}"


for particular solution:

"y_{p1}=Ax^2+Bx+C"

"2A-3(2Ax+B)+2(Ax^2+Bx+C)=x^2"

"A=1\/2"


"2A-3B+2C=0"

"-6A+2B=0\implies B=3/2"

"C=(3B-2A)\/2=(9\/2-1)\/2=7\/4"


"y_{p2}=Acosx+Bsinx"

"-Acosx-Bsinx-3(-Asinx+Bcosx)+2(Acosx+Bsinx)=sinx"

"A-4B=0"

"3A+B=1"

"B=1\/13,A=4\/13"


"y=y_c+y_{p1}+y_{p2}=c_1e^x+c_2e^{2x}+x^2\/2+3x\/2+7\/4+4cosx\/13+sinx\/13"